OpenStudy (anonymous):

The centripetal force on a 0.82kg object on the end of a 2.0m massless string being swung in a horizontal circle is 4.0N. What is the tangential velocity of the object?

OpenStudy (amilapsn):

What is the relationship between 4N, r & V?

OpenStudy (anonymous):

it makes an angle?:o

OpenStudy (amilapsn):

What?

OpenStudy (amilapsn):

@myphuong are you there?

OpenStudy (mrnood):

The centripetal force = mv^2/r

OpenStudy (nincompoop):

\[v_T = \frac{ds}{dt}=\frac{dr \theta}{dt} \rightarrow r \frac{d \theta}{dt}=r \omega\] http://www.dummies.com/how-to/content/calculating-tangential-velocity-on-a-curve.html

OpenStudy (anonymous):

I'm late, but thanks guys (:

OpenStudy (mrnood):

What else do you want?: It says a horizontal circle It gives you m and r and F I gave you F=mv^2/r so calculate V